c) The following data relates to the market value of economic transactions for the three main sectors of your country’s economy.

SECTOR
VALUE OF OUTPUT
PURCHASES FROM OTHER FIRMS
Agriculture
300
160
Manufacturing
200
150
Services
150
120

Required:

(i) Compute the gross domestic product (GDP) and explain the method used and why. (2 Marks)



(ii) Given that: depreciation = 90, indirect taxes = 70, subsidies = 30, payments to factors of production from abroad = 20, payments to foreign factors = 40.

Compute:


1) Gross national product (GNP) at market prices. (1 Mark)


2) Net national product (NNP) at market prices. (1 Mark)




3) Net national product (NNP) at factor cost. (1 Mark)


4) Net domestic product (NDP) at factor cost . (1 Mark)
1
Expert's answer
2020-07-06T17:34:39-0400

i) Value Added Method (Production)

To calculate GDP, only final goods and services are taken in order to exclude a double in the calculations and overstatement of GDP. Intermediate goods, which are the raw materials for the production of the final product, are not taken into account.


"(300-160)+(200-150)+(150-120)=220"

ii)

1)

"220+20-40=200"

2)

"200-90=110"

3)

"200-70+30=160"

4)

"220-90=130"